If 7.0 moles of nh3 reacts with 14 moles of o2, how many moles of no2 will be produced?

Calculate how many moles of excess reactant will remain
1
Expert's answer
2020-04-13T02:26:56-0400

4NH3 + 7O2 = 4NO2 + 6H2O

According to the equation NH3 reacts completely, O2 is in the excess.

7 mol of NO2 will be produced.

According to the equation 7mol*7/4 = 12. 25 mol of O2 reacts.

14 - 12.25 = 1.75 mol of O2 will remain.